Your name: Your Team Letter:

Group member’s name:

Group Citizenship Assignment

Choose the category that best fits this person’s behavior:

• Is this person timely in completing their portion of group assignments?

Does not complete Never finishes Sometimes finishes Always finishes
assignments  assignments on time assignments on time assignments on time
0  1  2  3

• Is this person contributing a fair amount to the project?

Does not contribute Contributes less Contributes what Contributes more
than is fair  is fair  than is fair on time
0  1  2  3

• Is this person prepared for meetings, including arriving on time?

Does not come Comes, but is never Is sometimes on time Is always on time
to meetings prepared or is and prepared and prepared on time

usually late

0  1  2  3

• Overall, to what extent is this perosn fullfilling your group’s Team Working Agreement

(contract)?

Does not meet Meets some of Meets most of  Meets or exceeds all
TWA obligations the obligations the obligations of the obligations on time
0  1  2  3

What is the one thing this person currently does that enhances the success of the group project?

What is the one thing this person could do that he/she is currently not doing, to enhance the success of the project?